how much pure acid should be mixed with 6 gallons of a 50% acid solution to get a 70% acid solution?
Answer by stanbon(75887)

how much pure acid should be mixed with 6 gallons of a 50% acid solution to get a 70% acid solution?
-------
Eqquation:
acid + acid = acid
x + 0.50(6) = 0.70(x+6)
---
Multiply thru by 100 to get:
100x + 50*6 = 70x + 70*6
30x = 20*6
x = 4 gallons (amt. of pure acid needed)
======================
Cheers,
Stan H.
=============